# ExPath: Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

### v1

@@ 3.3.2. Pathway-wise global aggregation
(6) $\mathcal{Q}=\left\{\bf{q}^{i}\mid\bf{q}^{i}\sim\text{Pathway}(v_{i},L),\,|\bf{q}^{i}|\leq L\right\}_{i=1}^{N},$
where $N$ is the number of nodes in the graph, and $\bf{q}^{i}$ represents the sampled pathway for node $v_{i}$. Each pathway $\bf{q}^{i}$ is a sequence of nodes $\{v_{i},v_{i_{1}},v_{i_{2}},\dots,v_{i_{L}}\}$, sampled according to a random walk process (Tonshoff et al., 2023). The sampling process $\text{Pathway}(v_{i},L)$ involves selecting a sequence of connected nodes starting from $v_{i}$. The selection of each subsequent node is determined probabilistically, guided by the graph adjacency structure.
Sequential Pathway Modeling. The forward propagation of the Mamba layer aggregates long-range dependencies along the sampled pathways. For each sampled pathway $\bf{q}^{i}\in\mathcal{Q}(X^{l})$, the Mamba layer processes the pathway sequentially as:
(7) $\Delta_{t}$ $=\tau_{\Delta}(f_{\Delta}(\mathbf{x}_{t}^{l})),\quad\mathbf{B}_{t}=f_{B}(\mathbf{x}_{t}^{l}),\quad\mathbf{C}_{t}=f_{C}(\mathbf{x}_{t}^{l}),$
$\mathbf{h}_{t}^{l}$ $=(1-\Delta_{t}\cdot\mathbf{D})\mathbf{h}_{t-1}^{l}+\Delta_{t}\cdot\mathbf{B}_{t}\mathbf{x}_{t}^{l}$
$X^{l+1}_{G}$ $=C\cdot h^{l+1}_{L},$
@@
3.3.3. Graph Classification
The updated node features are aggregated using a max pooling to generate a graph representation. This representation is passed through an MLP layer for classification:
(8) $y=\text{MLP}\big{(}\text{MaxPooling}\big{(}\{h_{v_{i}}\}_{i=1}^{N}\big{)}\big{)},$
where $N$ is the number of nodes, and $y$ is the predicted class label for the pathway. The model is trained using the cross-entropy loss: $\mathcal{L}_{\text{cross-entropy}}=-\sum_{i=1}^{C}y_{i}\log\hat{y}_{i},$ where $C$ is the number of classes, $y_{i}$ is the ground truth label, and $\hat{y}_{i}$ is the prediction.
3.4. PathExplainer: Targeted Pathway Inference
PathExplainer directly infer subgraphs to generate targeted pathways by leveraging the interpretability of PathMamba. Vallina GNNexplainers (Ying et al., 2019; Luo et al., 2020), which focus primarily on the node or edge level, often struggle to capture the global structures at the pathway level. In contrast, PathExplainer introduces a key technical novelty by training pathway masks, where entire pathways (i.e., sequences of connected nodes and edges) are selectively masked during training to evaluate their contributions to PathMamba.
PathExplainer formalizes the identification of important subgraphs as an optimization problem. For a given graph $G=(V,E)$, the explanation is defined as $(G_{S},F_{S})$, where $G_{S}\subseteq G$ is the subgraph and $F_{S}$ represents the selected features. The explanation is derived by optimizing the mutual information $\mathcal{MI}(\cdot)$ between the subgraph and the model’s prediction:
(9) $\max_{G_{S},F}\mathcal{MI}(Y,(G_{S},F))=H(Y)-H(Y\mid G=G_{S},X=F_{S}),$
where $H(Y)$ is the entropy of the predictions and $H(Y\mid G=G_{S},X=F_{S})$ is the conditional entropy given the explanation.
The optimization is approached by learning a pathway mask $M$ for the sampled pathway’s edges and nodes. To enhance the interpretability and biological relevance of the pathway mask, random pathways $\mathcal{Q}$ are sampled as described in Section 3.3.2. For each node $v_{i}$, a single random pathway $q_{i}$ of length up to $L$ is sampled. These pathways are then used to restrict the mask learning process to edges within the sampled pathways, ensuring that the learned $M$ focuses on them. Specifically, the adjacency matrix $A$ is modified based on the pathway mask $M$ as $A^{\prime}=A\odot\sigma(M)$, where $\sigma$ denotes the sigmoid function. Similarly, the features are masked as $X^{\prime}=X\odot\sigma(M)$. The loss function for PathExplainer combines two components: a cross-entropy term for prediction consistency and regularization terms for sparsity:
$\mathcal{L}_{\text{mask}}:=-\sum_{c=1}^{C}1[y=c]\log P_{\Phi}(Y=c\mid G=A^{\prime},X=X^{\prime})d+\lambda\|M\|\end{aligned}$
where $\|M\|$ encourages sparsity in the edge selection, and $\lambda$ balances the trade-off between the classification loss and the sparsity regularization. Hence, the identified important subgraphs and node features (referring to AA sequence data) that contribute most to specific bio-networks can considered as targeted pathways.
4. Experiments and Results
Table 1. Baseline comparison results on bio-network classification. The best-performing and second-best results are highlighted in bold and underline, respectively. The gray-shaded rows indicate PathMamba (650M) with different ESM-2 parameter settings.
@@ 4. Experiments and Results
Dataset and Preprocessing. We collected all available human pathway networks from the widely used knowledge database, KEGG (Kanehisa and Goto, 2000). Our dataset consists of four main classes: Human Diseases, Metabolism, Molecular and Cellular Processes, and Organismal Systems (Kanehisa et al., 2024), covering 301 bio-networks. We searched for and downloaded all the raw data for the human pathway network (referred to as Homo sapiens) using KEGG APIs. To construct high-quality, trainable graphs, for nodes, we ensured that all protein nodes in the network were linked to their reference AA sequence data (Kanehisa et al., 2016). Protein nodes with a lack of or incomplete sequencing data were excluded. For edges, we streamlined the network structure by removing redundant or biologically insignificant interactions. We transformed preprocessed data into machine-learning-ready revision. The detailed data description and preprocessing pipeline can be found in Appendix C and Table 4.
@@
Experimental Setup. We conducted 10-fold stratified K-Fold cross-validation repeated 5 times. The mean and standard deviation of the results across all folds were reported for evaluation. The hyperparameters were determined using grid search to identify the optimal configuration for the model. Training for all models was accomplished on NVIDIA A6000 GPU and Xeon Gold 6258R CPU.
4.1. Experiment-I: Pathway Representation
Objective. Evaluate the classification performance on unseen bio-networks, in line with Hypothesis-1, and benchmark the results against baseline models.
Baselines and Metrics. We collected baselines from both message-passing GNNs and more advanced graph models, including GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), GPS (Rampášek et al., 2022), and Graph-Mamba (Wang et al., 2024). The detailed introduction of these baselines and the selection motivation can be found in Appendix D. We used precision, recall, and overall accuracy for the performance evaluation. We used 650M ESM-2 for PathMamba and all baselines.
Results. Table 1 presents PathMamba achieves the highest accuracy (0.754), outperforming all GNNs, GPS (0.726), GraphMamba (0.723). Furthermore, it secures best or second-best positions across all functional categories, demonstrating its robust ability to generalize across diverse pathway structures. The superior performance of the proposed method highlights its effectiveness in extracting and leveraging biologically meaningful structural information from pathway networks. The gray-shaded rows indicate the results of removing ESM-2 and modifying the model size in terms of F1 scores. When ESM-2 is removed, the accuracy deteriorate significantly (0.75 → 0.44), highlighting the importance of AA-seq and the limitations of prior studies that were unable to leverage this information. Increasing the model size gradually improves the results; however, accuracy does not incline with the 3B model (0.754 → 0.742). This suggests that excessively large features may lead to overfitting or noise, particularly in capturing functional pathways. Table 2 compares the training and inference times of our model with other expressive hybrid models, using a batch size of 32. Our training time is 30% faster than GPS and inference time is 27% faster than Graph-Mamba (complexity analysis can be found in Appendix B).
@@ 4.1. Experiment-I: Pathway Representation
Table 2. The computational efficiency comparison with hybrid models, including both training and inference runtime.
@@
4.2. Experiment-II: Pathway Inference
Objective. Quantify the fidelity of extracted subgraphs, following our Hypothesis-2, and validate the importance of pathways specific to biological functions.
Baselines and Metrics. We collected baselines from conventional statistical methods (Random Sampling (Kashtan et al., 2004b), PersonalizedPageRank (Iván and Grolmusz, 2011), and MinimumDominatingSet (Nacher and Akutsu, 2016; Wuchty, 2014)), gradient-based methods (Saliency (Simonyan et al., 2013), InputXGradient (Shrikumar et al., 2017), Deconvolution (Mahendran and Vedaldi, 2016), ShapleyValueSampling (Strumbelj and Kononenko, 2010), and GuidedBackpropagation (Springenberg et al., 2014)), and GNN-specific explainer methods (GNNExplainer (Ying et al., 2019) and PGExplainer (Luo et al., 2020)). All details can be found in Appendix D. We evaluated the distinctiveness of the pathways inferred by PathExplainer using fidelity metrics, specifically Fidelity+ and Fidelity-. Fidelity+ measures how well the important features identified by the model contribute to accurate predictions. In contrast, the fidelity- evaluates the drop in prediction accuracy when the identified important features are retained while others are removed. All details can be found in Appendix D and E.
@@ 4.2. Experiment-II: Pathway Inference
Figure 3. The fidelity+ (necessity ↑) and fidelity- (sufficiency ↓) scores of extracted subgraphs. Our PathExplainer achieves the best performance with the lowest fidelity+ and fidelity- scores, indicating its superior ability to produce robust and meaningful pathway networks.
Figure 4. UpSet plot of enriched GO terms across four pathway classes, based on top feature sets from subgraphs for different methods. Orange indicates GO terms uniquely enriched in one class, and blue represents GO terms shared across multiple classes. RSS, MDS, and PPR stand for Random Sampling, Minimum Dominating Set, and Personalized PageRank, respectively.
Results. Figure 3 illustrates that PathExplainer achieves the highest fidelity+ and the lowest fidelity-, indicating its ability to explain necessary and sufficient subgraphs effectively. GNN-specific or gradient-based methods (blue points) show lower fidelity- compared to traditional methods (green points), demonstrating that the learned AA-seq enables the identification of sufficient subgraphs. PGExplainer exhibits lower fidelity+ than GNNExplainer, suggesting that instance-based approaches may be better suited for explaining diverse pathways.
4.3. Experiment-III: Biological Meaningfulness
@@
Overview. We proposed an evaluation workflow to analyze the biological significance of the extracted subgraphs and pathways. This workflow directly integrates the weighting/ranking scores of pathway inferred by PathExplainer into biological metrics, enabling the direct quantification of outputs from deep learning models. Specifically, this includes: Breadth: The diversity of biological functions represented by the subgraphs. Depth: The extent to which gene nodes contribute to these biological functions. Reliability: The robustness and statistical significance of the analyses targeting these biological functions.
@@
Setup and Metrics. To this end, we designed experiments centered on Gene Ontology (GO) analysis (Ashburner et al., 2000), focusing on the nodes within the extracted subgraphs. The results provide a list of GO terms highlighting the biological functions most significantly represented in the input gene (corresponding to protein) nodes (Ashburner et al., 2000). Breadth was assessed using the Number of Enriched Biological Functions (#EBF). A higher #EBF indicates broader functional diversity within the subgraph. Depth was evaluated using a new metric called the Enrichment Contribution Score (ECS). ECS evaluates the relative contribution of the top-weighted genes, denoted as $G_{\text{Top}}$. Here we directly framed the inferred pathways as the targets of ECS to generate $G_{\text{Top}}$ for evaluations by following steps: Let $G=\{g_{1},g_{2},\dots,g_{n}\}$ represent the ranked list of gene nodes, sorted by the importance weights of pathways (inferred directly by PathExplainer) $w=\{w_{1},w_{2},\dots,w_{n}\}$, where $w_{1}\geq w_{2}\geq\dots\geq w_{n}$. Define $G_{\text{Top}}=\{g_{1},g_{2},\dots,g_{\text{Top}}\}$ to include only genes with the top weights, selected based on a ratio $R\%$ (defaulted as 30%), as a subset of $G$. Then, perform GO analysis based on $G_{\text{Top}}$ for each input subgraph $S_{i}$. The ECS is calculated as the average number of enriched items for each gene in $G_{\text{Top}}$ across all subgraphs $S_{i}$, and is defined as:
where $P$ is the total number of tested subgraphs, $GO_{\text{top-enriched}}(S_{i})$ is the set of enriched GO terms for subgraph $S_{i}$ based on $G_{\text{Top}}$, and $|G_{\text{Top}}|$ is the number of genes in the subset $G_{\text{Top}}$. Moreover, Reliability assessed using P-value. We accepted the item only with a P-value lower than 0.05, the average P-value reported here is naturally lower than this threshold. A lower average indicates greater reliability in the enrichment results across the subgraphs. The details of GO and metrics can be found in Appendix E.
@@
Table 3. Biological meaningfulness comparison results on subgraphs extracted by different methods. The best-performing results are highlighted in bold. The second-best results are highlighted in underline.
@@
Figure 4 evaluates the differences in enriched GO terms across four pathway classes based on top gene sets from subgraphs extracted by different methods. Compared with other methods, the upset plot reveals that PathMamba-PathE identifies the most extensive sets of unique GO terms (shown as the orange bars and links) across all four pathway classes while maintaining fewer shared terms (shown as the blue bars and links) among different classes. This suggests that PathMamba-PathE tends to assign appropriate weights to genes based on their importance within the network and effectively captures the distinct biological roles of top-ranked genes in specific pathway classes.
4.4. Case Study: T Cell Receptor Signaling Pathway Evaluation
@@
Setup. The T cell receptor (TCR) signaling pathway is a classic example of well-characterized human pathways. It would be helpful if parts implying key aberrations could be found as sub-graphs in the whole pathway graph. In this case study, we compare subgraphs extracted by two methods: TCR Subgraph A, generated using the RSS method, and Subgraph B, obtained via our proposed method. Each method selects the top 10% highest-ranked nodes and their associated edges to construct a representative subgraph. The detailed information and results analysis of the case study can be found in Appendix G.
TCR Subgraph A: The RSS Method. As shown in Figure 5, subgraph A, generated by the RSS method, distributes high scores uniformly across a broad range of nodes within the TCR pathway. However, this hints at unnatural, fragmented signal propagation, as highlighted by numerous discrete red-marked nodes.
TCR Subgraph B: The Proposed Method. In contrast, as shown in Figure 5 subgraph B, the subgraph extracted by our method exhibits a strong focus on the PI3K-AKT signaling axis (Vara et al., 2004) and the downstream components of the NF-$\kappa$B (Dolcet et al., 2005) pathway, as highlighted by the coherent red-marked path.
@@
Discuss. In summary, the proposed method’s subgraphs align with real-world pathway analysis practices’ needs: maintaining signal continuity within regulatory cascades and even holding relatively long signaling paths, making it more suitable for focused analyses of bio-network regulatory mechanisms.
Figure 5. Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. TCR Subgraph A is from the RSS method, and TCR Subgraph B is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes.
5. Related work
Existing methods for analyzing the structure of biological networks typically represent these networks as graphs, aiming to infer subgraphs or extract relevant interactions, and can generally be categorized into statistical topology-driven and data-driven deep graph learning methods.
Topology-driven methods utilize statistical metrics on structural properties of graphs, such as node degrees (Karger, 1994; Kashtan et al., 2004a), centrality (Haynes et al., 2013; Wang et al., 2014; Nacher and Akutsu, 2016), betweenness, or PageRank scores (Iván and Grolmusz, 2011; Maehara et al., 2014) to infer which substructures or edges exert a more significant influence on the overall topology, thereby identifying more targeted interactions among genes or proteins.
Deep graph learning methods incorporate experimental data during the learning process by embedding data as node representations. They train graph neural networks (GNNs) with suitable objectives, such as link prediction or graph reconstruction (Yue et al., 2020; Zhang et al., 2021; Muzio et al., 2021), and the links that contribute most to these objectives can be considered the targeted interactions. For instance, the classic GCN algorithm, GraphSAGE (Hamilton et al., 2017), has been validated on protein-protein interaction (PPI) datasets to predict protein functions within networks. The work of (Gligorijević et al., 2021) introduced a GCN-based method for predicting protein functions, leveraging sequence features derived from a protein language model alongside structural information. Chen et al. (Cheng et al., 2021) employed a graph attention network to extract drug and protein AA-Seq features for predicting drug-target interactions. Moreover, GNN models have been applied to incorporate RNA-Seq data, for tasks like predicting disease states and cell-cell relationships (Wang et al., 2021; Ravindra et al., 2020).
6. Conclusion

### v2

@@ Introduction
$\mathcal{Q}=\left\{\bf{q}^{i}\mid\bf{q}^{i}\sim\text{Pathway}(v_{i},L),\,|\bf{q}^{i}|\leq L\right\}_{i=1}^{N},$ (6)
where $N$ is the number of nodes in the graph, and $\bf{q}^{i}$ represents the sampled pathway for node $v_{i}$. Each pathway $\bf{q}^{i}$ is a sequence of nodes $\{v_{i},v_{i_{1}},v_{i_{2}},\dots,v_{i_{L}}\}$, sampled according to a random walk process (Tonshoff et al. 2023). The sampling process $\text{Pathway}(v_{i},L)$ involves selecting a sequence of connected nodes starting from $v_{i}$. The selection of each subsequent node is determined probabilistically, guided by the graph’s adjacency structure.
- Sequential Pathway Modeling. The forward propagation of the Mamba layer aggregates long-range dependencies along the sampled pathways. The selective sequential modeling of Mamba is well-suited for capturing such path information. For each sampled pathway $\bf{q}^{i}\in\mathcal{Q}(X^{l})$, the Mamba layer processes the pathway sequentially as:
$\Delta_{t}$ $=\tau_{\Delta}(f_{\Delta}(\mathbf{x}_{t}^{l})),\quad\mathbf{B}_{t}=f_{B}(\mathbf{x}_{t}^{l}),\quad\mathbf{C}_{t}=f_{C}(\mathbf{x}_{t}^{l}),$ (7)
$\mathbf{h}_{t}^{l}$ $=(1-\Delta_{t}\cdot\mathbf{D})\mathbf{h}_{t-1}^{l}+\Delta_{t}\cdot\mathbf{B}_{t}\mathbf{x}_{t}^{l},\quad X^{l+1}_{G}=C\cdot h^{l+1}_{L},$
@@
Afterward, we apply max pooling over the node features, i.e., $\{h_{v_{i}}\}_{i=1}^{N}$, followed by an MLP and softmax activation for the classification task.
PathExplainer: Targeted Pathway Inference
PathExplainer directly infers subgraphs to generate targeted pathways by leveraging the interpretability of PathMamba. Vallina GNNexplainers (Ying et al. 2019; Luo et al. 2020), which focus primarily on the node or edge level, often struggle to capture the global structures at the pathway level. In contrast, PathExplainer introduces novel pathway mask training, where entire pathways (i.e., connected nodes and edges) are selectively masked during training to evaluate their contributions to PathMamba.
Theoretical Objective.
[…]
Optimization Framework.
The optimization is approached by learning a pathway mask $\mathcal{M}$ for the sampled pathway’s edges and nodes based on random pathways $\mathcal{Q}$ as described in Section Pathway-wise global aggregation.. For each node $v_{i}$, a random pathway $q_{i}$ of length up to $L$ is sampled. These pathways are then used to restrict the mask learning process within the sampled pathways, ensuring that the learnable pathway mask $\mathcal{M}$ focuses on them. Specifically, the targeted subgraph $\hat{\mathcal{G}}$ is inferred based on $\mathbf{M}$ as: $\hat{\mathcal{G}}=\mathcal{G}\odot\sigma(\mathcal{M})$, where $\sigma$ denotes the sigmoid function. The loss function for PathExplainer combines two components: a cross-entropy term for prediction consistency and regularization terms for sparsity:
[…]
where $\|M\|$ encourages sparsity in the edge selection, $Y$ is a random variable representing labels {1, 2, …, C}, and $\lambda$ balances the trade-off between the classification loss and the sparsity regularization. Hence, the identified important subgraphs and node features that contribute most to specific bio-networks are considered as targeted pathways.
Table 1: Baseline comparison results on bio-network classification. The best and second-best results are highlighted in bold and underline, respectively. The gray-shaded rows indicate PathMamba with different ESM-2 (encoder) parameter settings.
@@
Theoretical Analysis for Targeted Pathway Fidelity
In this section, we place our method within the Weisfeiler–Lehman (WL) hierarchy to characterize its expressive power (Details can be found in Appendix A). By proving that our explainer goes beyond the 1-WL limitation, we ensure that the extracted pathways capture higher-order structural patterns, establishing a theoretical upper bound on fidelity and inference ,supporting the empirical results.
{restatable}
lemmaexprforexpl(Expressiveness for explanations). When combined with higher expressive models (e.g., it distinguishes more graphs), PathExplainer can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs). In contrast, less expressive models merge different graphs into larger equivalence classes, leading to non-unique, less granular explanations.
In Appendix B, we prove this by showing that the expressiveness of the underlying graph classifier $f$ determines the granularity of equivalence classes, with more expressive models enabling finer distinctions between graphs.
Lemma 1 (Comparison with k-WL test).
Proof.
The proof of this theorem directly comes from the recent work (Tonshoff et al. 2023; Behrouz and Hashemi 2024). They prove a similar theorem using 1-d CNN (Tonshoff et al. 2023) or SSM (Behrouz and Hashemi 2024) with randomly sampled subgraph. ∎
Lemma 2 (Comparison with 1-WL test).
PathMamba is strictly more expressive than 1-WL GNNs.
Proof.
We first note that PathMamba contains the GIN as a sub-module, which has the same expressive power as the 1-WL test (Xu et al. 2019). Therefore, PathMamba is at least as expressive as 1-WL GNNs. By Lemma 1, there are graphs that cannot be distinguished by $1$-WL GNNs, but can be distinguished by PathMamba. Consequently, PathMamba is strictly more expressive than 1-WL GNNs. ∎
Theorem 1 (Explanations of ExPath).
Based on Lemma Theoretical Analysis for Targeted Pathway Fidelity, 1, and 2, ExPath can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs) than 1-WL GNN-based methods, and not bounded by any WL GNN methods.
Experiments and Results
@@
Dataset and Preprocessing. We collected all available human pathway networks from the widely used knowledge database, KEGG (Kanehisa and Goto 2000). Our dataset consists of four main classes: Human Diseases, Metabolism, Molecular and Cellular Processes, and Organismal Systems, covering 301 bio-networks. For nodes, we ensured that all protein nodes in the network were linked to their reference AA sequence data. The detailed data description and preprocessing can be found in Appendix D and Table 5.
@@
Experimental Setup. We conducted 10-fold stratified K-fold cross-validation repeated five times. The optimal hyperparameters were determined using grid search. Training for all models was implemented on NVIDIA A6000 GPU and Xeon Gold 6258R CPU.
@@
Table 2: The computational efficiency comparison with hybrid models, including both training and inference runtime.
@@
Experiment-I: Pathway Learning
Objective. This experiment aims to evaluate whether ExPath can classify diverse bio-networks and benchmark its performance against baseline models.
Baselines and Metrics. We collected baselines from both message-passing GNNs and more advanced graph models, including GCN (Kipf and Welling 2017), GraphSAGE (Hamilton, Ying, and Leskovec 2017), GAT (Veličković et al. 2018), GIN (Xu et al. 2019), GPS (Rampášek et al. 2022), and Graph-Mamba (Wang et al. 2024). The detailed introduction of these baselines and the selection motivation can be found in Appendix E. We employed precision, recall, and overall accuracy for the performance evaluation. We used 650M ESM-2 for PathMamba and all baselines as the node feature encoding model.
Results. Table 1 demonstrates that PathMamba achieves the highest accuracy (0.754), outperforming all GNNs, GPS (0.726), GraphMamba (0.723). Furthermore, it secures best or second-best positions across all functional categories, demonstrating its robust ability to generalize across diverse pathway structures. The gray-shaded rows indicate the results of removing ESM-2 and modifying the model size in terms of F1 scores. When ESM-2 is removed, the accuracy decreases significantly (0.75 → 0.44). The results highlight the importance of AA-seq and the limitations of prior studies that were unable to leverage this information.
Table 2 compares the training and inference times of our model with other expressive hybrid models, using a batch size of 32. Our training time is 30% faster than GPS, and inference time is 27% faster than Graph-Mamba (complexity analysis can be found in Appendix C).
@@
Figure 3: Fidelity+ (necessity $\uparrow$) and Fidelity- (sufficiency $\downarrow$) scores of extracted subgraphs. Our PathExplainer achieves the best performance on both metrics.
Experiment-II: Pathway Inference
Objective. This experiment aims to quantify the fidelity of extracted subgraphs using PathExplainer and validate the importance of pathways specific to biological functions.
Baselines. We collected baselines from three categories (i) conventional statistical methods: Random Sampling (Kashtan et al. 2004b), PersonalizedPageRank (Iván and Grolmusz 2011), and MinimumDominatingSet (Nacher and Akutsu 2016; Wuchty 2014)); (ii) gradient-based methods: Saliency (Simonyan, Vedaldi, and Zisserman 2013), InputXGradient (Shrikumar, Greenside, and Kundaje 2017), Deconvolution (Mahendran and Vedaldi 2016), ShapleyValueSampling (Strumbelj and Kononenko 2010), and GuidedBackpropagation (Springenberg et al. 2014); (iii) GNN-specific explainer method: GNNExplainer (Ying et al. 2019) and PGExplainer (Luo et al. 2020). All details can be found in Appendix E.
Metrics. We evaluated the distinctiveness of the pathways inferred by PathExplainer using fidelity metrics, Fidelity+ and Fidelity-. Fidelity+ measures how well the selected features support accurate predictions, while Fidelity- checks how much accuracy drops when only those features are kept. We further evaluated the length of pathways. Max Path Length captures the longest simple path in each subgraph, reflecting whether long signaling chains are retained. Average Diameter measures the typical node-to-node distance, showing how spread out the nodes remain after extraction. All details can be found in Appendix E and F.
Table 3: Comparison of pathway-preservation ability across subgraph extraction baselines. Higher path length and diameter indicate better retention of long-range interactions.
Methods Max Path Length Average Diameter
Minimum Dominating Set 6 2.00
Random Sampling 11 2.90
Personalized Page Rank 4 1.53
PathMamba-GNNExplainer 9 2.90
GPS-PathExplainer 12 3.95
ExPath (Ours) 16 4.20
Results. Figure 3 shows that PathExplainer achieves the highest fidelity+ and the lowest fidelity-. The main reason is that GNNExplainers optimise a mask for each individual node or edge level, whereas PathExplainer infers pathways (subgraphs) as a single coherent unit. Deconvolution simply aggregates all edges with large gradients, it almost covers every active edge. This pushes low fidelity-, but since it retains many redundant edges, removing them hardly changes the output, so fidelity+ (necessity) stays low. GNN-specific or gradient-based methods (blue points) show lower fidelity- compared to traditional methods (green points), indicating that the learned AA-seq enables the identification of sufficient subgraphs.
Table 3 presents that our method attains up to 4 × longer preserved paths and up to 2.7 × larger diameters than competing approaches. This supports that the identified sufficient and necessary features capture biologically meaningful pathways and meets our Expectation.
Experiment-III: Biological Meaningfulness
@@
Objective. We propose an evaluation workflow to analyze the biological significance of the subgraphs and pathways extracted from our method. This workflow should integrate the weighting/ranking scores of pathway inferred by ExPath into biological metrics, enabling the direct quantification of outputs from the models.
@@
Proposed Evaluation Metrics. We designed experiments centered on Gene Ontology (GO) analysis (Ashburner et al. 2000), focusing on the nodes within the extracted subgraphs. The results provide a list of GO terms highlighting the biological functions most significantly represented in the input gene (corresponding to protein) nodes (Ashburner et al. 2000). Then we proposed Number of Enriched Biological Functions (#EBF) and Enrichment Contribution Score (ECS) to evaluate breadth and depth of the extracted functions (Yang et al. 2024). A higher #EBF indicates broader functional diversity within the subgraph. ECS evaluates the relative contribution of the top-weighted genes. The details and definitions of metrics can be found in Appendix F.
@@
Table 4: Biological meaningfulness comparison results: The best-performing results are highlighted in bold. The second-best results are highlighted in underline.
@@
Figure 4: UpSet plot of enriched GO terms across four pathway classes, based on top feature sets from subgraphs for different methods. Orange indicates GO terms uniquely enriched in one class, and blue represents GO terms shared across multiple classes. RSS, MDS, and PPR stand for Random Sampling, Minimum Dominating Set, and Personalized PageRank, respectively.
Figure 5: Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. TCR Subgraph A is from the RSS method, and TCR Subgraph B is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes. A full pathway graph can be found in Appendix G.
@@
Figure 4 evaluates the differences in enriched GO terms across four pathway classes based on top gene sets from subgraphs extracted by different methods. The upset plot reveals that PathMamba-PathE identifies the most extensive sets of unique GO terms (shown as the orange bars and links) across all four pathway classes while maintaining fewer shared terms (shown as the blue bars and links) among different classes. This suggests that PathMamba-PathE tends to assign appropriate weights to genes based on their importance within the network and effectively captures the distinct biological roles of top-ranked genes in specific pathway classes.
Experiment IV: PoC of Biological Case Study
@@
Setup. We make a Proof-of-Concept case analysis using the T cell receptor (TCR) signaling pathway, which is a well-characterized human pathway. In this case study, we compare subgraphs extracted by two methods: TCR Subgraph A, generated using the RSS method, and Subgraph B, obtained via our proposed method. Each method selects the top 10% highest-ranked nodes and their associated edges to construct a representative subgraph. The detailed information and analysis of the case study can be found in Appendix H.
Results. In Figure 5, the upper subfigure presents the TCR subgraph extracted by the RSS method, while the lower one corresponds to our ExPath. In subgraph A, generated by the RSS method, high scores are distributed uniformly across a broad range of nodes within the TCR pathway. However, this suggests unnatural, fragmented signal propagation, as evidenced by the numerous isolated red-marked nodes and a green broken connection in PDPK1. In contrast, subgraph B, extracted by our method, exhibits a strong focus on the PI3K-AKT signaling axis (Vara et al. 2004) and the downstream components of the MAP3K8 survival (Dolcet et al. 2005), as highlighted by a coherent red-marked path.
@@
Discussion. In summary, the extracted subgraphs by ExPath align with the needs of real-world pathway analysis practices: maintaining signal continuity within regulatory cascades and even accommodating relatively long signaling paths, making them more suitable for focused analyses of bio-network regulatory mechanisms.
Conclusion